# A data-driven method to estimate contamination from light ion beam transmutation at colliders

### v1

@@ IV Complicating factors
Another method for dealing with pileup effects involves careful planning of the beam conditions when collecting the dataset to be analyzed. Modern colliders are able to add a small spatial offset to colliding particle beams, which effectively reduces the instantaneous luminosity delivered to an interaction point. This beam separation process can be used manipulate the luminosity profile (and therefore the pileup profile) seen by an experiment. In particular, the instantaneous luminosity can be ‘leveled’ by separating the beams to achieve a desired pileup value, and then slowly bringing the beams back together over time to compensate for decreasing beam intensities. Requesting leveling at a fixed value for the duration of a fill is a natural way that an experiment can remove the time-dependent pileup effects that could confound contamination extraction.
@@
The leveling process does come with a downside: the separation of the beams reduces the integrated luminosity sampled by an experiment, potentially hindering luminosity-hungry physics programs that are reliant on very rare processes. It is worth noting that a number of ‘hybrid’ leveling strategies could be pursued to reach a reasonable compromise between the sampled integrated luminosity and the level of precision that beam contamination can be extracted. A schematic sketch of these strategies, shown as a plot of the collected instantaneous luminosity vs time, is shown in Fig. 7. The most conservative strategy is to fully level for the duration of the fill, which effectively removes the pileup uncertainty from this method. A more aggressive strategy would be to request leveling at the beginning of a fill to establish the reference control region and then periodically return to the leveled value to establish measurements of the contamination versus time. These measurements can then be interpolated to estimate the full time-dependence of the contamination effect for the entire duration of the fill. Alternatively, just the beginning and the end of the fill can be leveled to establish a measurement of the contamination at the end of the fill, where it is expected to be largest. This strategy could be appropriate if only an upper limit on the amount of contamination is needed. Perhaps the most aggressive leveling strategy would be leveling at the beginning of the fill to a value that the luminosity is expected to naturally decay to at the end of the fill. This would be essentially equivalent to leveling at both the beginning and the end of the fill without actually realizing any luminosity losses at the end of the fill.
@@
Multiple contaminants. For light ion beams having $Z/A=0.5$, many potential contaminant ions are able to continue to circulate in the collider. To illustrate our method of estimating transmutation backgrounds, we made a simplifying assumption that there is one dominant contaminant species ($^{4}\text{He}$ in the case of $^{16}\text{O}$ beams). The presence of more than one contaminant species potentially complicates the structure of the contaminant $N_{\text{trk}}$ distribution, as it is would no longer be expected to have only one clear “knee” above which the distribution quickly falls to zero. The method we propose does not depend on the specific shape of the overall contaminant $N_{\text{trk}}$ distribution (and in fact it could be used to extract the shape of this distribution as shown in the bottom panel of Fig. 6) as long as the value of $N_{\text{trk}}^{\text{cut}}$ is selected such that all contaminant events produce less multiplicity than the chosen threshold. However, additional tails coming from heavier contaminant collisions, such as $^{14}\text{N}^{16}\text{O}$ in the case of $^{16}\text{O}$ beams will have a knee structure that is very close to the upper $N_{\text{trk}}$ limit for $^{16}\text{O}^{16}\text{O}$, and therefore could potentially bleed into the high purity control region if the $N_{\text{trk}}^{\text{cut}}$ value is too low. To evaluate the impact of this potential effect, an upward scan of the experimenter’s nominal $N_{\text{trk}}^{\text{cut}}$ choice can be performed to search for additional “knee” structures in the extracted contaminant distribution. A lack of observable structures above the nominal $N_{\text{trk}}^{\text{cut}}$ choice would imply that all of the relevant contaminants are accounted for by the current choice, whereas the presence of additional structures would suggest that the $N_{\text{trk}}^{\text{cut}}$ value needs to be raised to account for additional contaminants. In this way, the experimenter can systematically explore the degree to which additional contaminants may be relevant. Statistical considerations for the event yield in the high purity control region are expected to limit the maximal value that $N_{\mathrm{trk}}^{\mathrm{cut}}$ can take, so experiments gathering a high rate of data may be able to be more conservative with their choice of $N_{\mathrm{trk}}^{\mathrm{cut}}$.
@@
Delayed start of data collection. Even if an ion beam starts its lifetime without any contamination, transmutation will begin essentially as soon as a sufficient instantaneous luminosity is achieved to catalyze a high rate of electromagnetic dissociation. Experiments are commonly unable to begin data collection immediately after this point because of detector protection requirements—some short period of time (typically less than a few minutes) is usually needed to ramp detector voltages, move non-static detector components, etc., to their nominal data-taking parameters. The result of this delay is that the reference control region will never be truly free of contamination. The presence of some contamination in this control region will result in an over subtraction in the region of interest, producing an underestimate of the total contamination rate.
@@
The effects of a delayed start of data collection can be estimated by repeating the contamination estimation with several different choices for the reference control region, where each additional choice is delayed by some some small positive time shift $\delta$. One can then plot the extracted contamination value in the region of interest versus $\delta$ and then extrapolate to what the true contamination value will be if $\delta=-\Delta$, where $\Delta$ is the time between the start of high-intensity collisions and the start of data collection. The functional form used in the extrapolation is not guaranteed to be linear, and the experimenter should determine what is a reasonable hypothesis based on the data or accelerator simulations.
@@ V Conclusion
In summary, a data-driven method to estimate potential beam transmutation contamination rates in light-ion collisions by exploiting their distinct event-occupancy distribution shapes and the time-dependence of expected transmutation effects has been proposed. Using HG-Pythia to generate OO and HeO events, a toy model for the total $N_{\mathrm{trk}}$ distribution as a function of time for a typical LHC fill was created. We demonstrate that the method is able to reliably estimate the level of beam contamination under some simplifying assumptions. Potential experimental difficulties with the proposed method are discussed and recommendations for mitigating these challenges are given. This study is intended to inform discussion of beam transmutation contamination and relevant analysis strategies in recently collected OO and NeNe datasets, as well as in future light-ion runs at the LHC and other ion-ion colliders. Finally, it is worth noting that a strong understanding of contamination effects may ultimately result in the ability to use contaminant collisions for novel physics analyses, such as the those proposed in Nijs and van der Schee (2025), but we leave those discussions for future work.

### v2

@@ IV Complicating factors
Another method for dealing with pileup effects involves careful planning of the beam conditions when collecting the dataset to be analyzed. Modern colliders are able to add a small spatial offset to colliding particle beams, which effectively reduces the instantaneous luminosity delivered to an interaction point. This beam separation process can be used manipulate the luminosity profile (and therefore the pileup profile) seen by an experiment. In particular, the instantaneous luminosity can be ‘leveled’ by separating the beams to achieve a desired pileup value and then slowly bringing the beams back together over time to compensate for decreasing beam intensities. Requesting leveling at a fixed value for the duration of a fill is a natural way that an experiment can remove the time-dependent pileup effects that could confound contamination extraction.
@@
The leveling process does come with a downside: the separation of the beams reduces the integrated luminosity sampled by an experiment, potentially hindering luminosity-hungry physics programs that are reliant on very rare processes. It is worth noting that a number of ‘hybrid’ leveling strategies could be pursued to reach a reasonable compromise between the sampled integrated luminosity and the level of precision that beam contamination can be extracted. A schematic sketch of these strategies, shown as a plot of the collected instantaneous luminosity vs time, is shown in Figure 7. The most conservative strategy is to fully level for the duration of the fill, which effectively removes the pileup uncertainty from this method. A more aggressive strategy would be to request leveling at the beginning of a fill to establish the reference control region and then periodically return to the leveled value to establish measurements of the contamination versus time. These measurements can then be interpolated to estimate the full time-dependence of the contamination effect for the entire duration of the fill. Alternatively, just the beginning and the end of the fill can be leveled to establish a measurement of the contamination at the end of the fill, where it is expected to be largest. This strategy could be appropriate if only an upper limit on the amount of contamination is needed. Perhaps the most aggressive leveling strategy would be leveling at the beginning of the fill to a value that the luminosity is expected to naturally decay to at the end of the fill. This would be essentially equivalent to leveling at both the beginning and the end of the fill without actually realizing any luminosity losses at the end of the fill.
@@
Multiple contaminants. For light ion beams having $Z/A=0.5$, many potential contaminant ions are able to continue to circulate in the collider. To illustrate our method of estimating transmutation backgrounds, we made a simplifying assumption that there is one dominant contaminant species ($^{4}\text{He}$ in the case of $^{16}\text{O}$ beams). The presence of more than one contaminant species potentially complicates the structure of the contaminant $N_{\text{trk}}$ distribution, as it is would no longer be expected to have only one clear “knee” above which the distribution quickly falls to zero. The method we propose does not depend on the specific shape of the overall contaminant $N_{\text{trk}}$ distribution (and in fact it could be used to extract the shape of this distribution as shown in the bottom panel of Fig. 6) as long as the value of $N_{\text{trk}}^{\text{cut}}$ is selected such that all contaminant events produce less tracks than the chosen threshold. However, additional tails coming from heavier contaminant collisions, such as $^{14}\text{N}^{16}\text{O}$ in the case of $^{16}\text{O}$ beams will have a knee structure that is very close to the upper $N_{\text{trk}}$ limit for $^{16}\text{O}^{16}\text{O}$, and therefore could potentially bleed into the high purity control region if the $N_{\text{trk}}^{\text{cut}}$ value is too low. To evaluate the impact of this potential effect, an upward scan of the experimenter’s nominal $N_{\text{trk}}^{\text{cut}}$ choice can be performed to search for additional “knee” structures in the extracted contaminant distribution. A lack of observable structures above the nominal $N_{\text{trk}}^{\text{cut}}$ choice would imply that all of the relevant contaminants are accounted for by the current choice, whereas the presence of additional structures would suggest that the $N_{\text{trk}}^{\text{cut}}$ value needs to be raised to account for additional contaminants. In this way, the experimenter can systematically explore the degree to which additional contaminants may be relevant. Statistical considerations for the event yield in the high purity control region are expected to limit the maximal value that $N_{\mathrm{trk}}^{\mathrm{cut}}$ can take, so experiments gathering a high rate of data may be able to be more conservative with their choice of $N_{\mathrm{trk}}^{\mathrm{cut}}$.
Figure 8: Percentage-level contamination at 12 hours as a function of the delayed time. The vertical black dashed line indicates the start of data collection at $t=0.5~\text{h}$. The solid red line shows a linear fit to the data points with $t\geq 0.5~\text{h}$ (solid black markers), while the open markers denote data points at $t<0.5$ h that are not included in the fit. The horizontal red dotted line denotes the “true” contamination at $t=0$, which is within 0.01% of the contamination extracted in the limit $t\to 0$, e.g., at $t=0.01$ h.
@@
Delayed start of data collection. Even if an ion beam starts its lifetime without any contamination, transmutation will begin essentially as soon as a sufficient instantaneous luminosity is achieved to catalyze a high rate of electromagnetic dissociation. Experiments are commonly unable to begin data collection immediately after this point because of detector protection requirements—some short period of time (typically less than a few minutes) is usually needed to ramp detector voltages, move non-static detector components to their nominal data-taking parameters, etc.. The result of this delay is that the reference control region will never be truly free of contamination. The presence of some contamination in this control region will result in an over subtraction in the region of interest, producing an underestimate of the total contamination rate.
@@
The effects of a delayed start of data collection can be estimated by repeating the contamination extraction using several choices of the reference control region, each corresponding to a different delay after the onset of high-intensity collisions. The extracted contamination at a fixed time (e.g., 12 hours) can then be plotted as a function of the delay to study how it varies with the start time. In Figure 8, the vertical dashed line at $t=0.5$ h marks the actual start of data collection. A linear fit is applied to the data for $t\geq 0.5$ h and extrapolated to earlier times to infer the true contamination at the onset of collisions. The extrapolation function is not guaranteed to be linear, and the experimenter should choose a reasonable functional form based on the data or accelerator simulations. The extrapolated value at an earlier time (eg: $t=0.01$ h) is compared to the true contamination calculated from Fig. 2, showing excellent agreement within 0.01%.
@@ V Conclusion
In summary, a data-driven method to estimate potential beam transmutation contamination rates in light-ion collisions by exploiting their distinct event-occupancy distribution shapes and the time-dependence of expected transmutation effects has been proposed. Using HG-Pythia to generate OO and HeO events, a toy model for the total $N_{\mathrm{trk}}$ distribution as a function of time for a typical LHC fill was created. We demonstrate that the method is able to reliably estimate the level of beam contamination under some simplifying assumptions. Potential experimental difficulties with the proposed method are discussed and recommendations for mitigating these challenges are given. This study is intended to inform discussion of beam transmutation contamination and relevant analysis strategies in recently collected OO and NeNe datasets, as well as in future light-ion runs at the LHC and other ion-ion colliders. Finally, it is worth noting that a strong understanding of contamination effects may ultimately result in the ability to use contaminant collisions for novel physics analyses, such as the those proposed in Ref. Nijs:2025qxm, but we leave those discussions for future work.